McCann Catching, Inc. has 2.00 million shares of stock outstanding. The stock currently sells for $12.75 per share. The firm’s debt is publicly traded and was recently quoted at 88.00% of face value. It has a total face value of $11.00 million, and it is currently priced to yield 9.00%. The risk free rate is 4.00% and the market risk premium is 8.00%. You’ve estimated that the firm has a beta of 1.20. The corporate tax rate is 36.00%.

The firm is considering a $41.97 million expansion of their production facility. The project has the same risk as the firm overall and will earn $10.00 million per year for 6.00 years.

What is the WACC for McCann Catching, Inc.?
Answer #1

cost of equity=4.00%+1.20*8.00%=13.60%

value of equity=2*12.75=25.50 million

value of debt=11.00*88.00%=9.68 million

WACC=(9.68/(9.68+25.50))*9.00%*(1-36.00%)+(25.50/(9.68+25.50))*13.60%

=11.44%

the above is answer..